A population of squirrels have 250 red-sided individuals and 750 tan-sided individuals. Assume that red is totally recessive. Calculate the allele frequencies of each allele.
1
Expert's answer
2018-05-15T06:15:16-0400
Red-sides animals result from occasional combination between two recessive alleles. This even happens with the probability, which is equal to squared frequency of the recessive allele in population.
The frequency of red-sides animals is equal to their probability, and for this population comprises 250 / (250+750) = 0.25.
Let q is the probability of recessive allele. Then, q^2 = 0.25, and q =0.5
Let p is the probability of dominant allele. Then, p=1-q, p=0.5